2 60 E 40 30ト . , , , D: 20 L 10 20 30 40 50 70 Tablet Devices (millions per year)

Identify the opportunity cost of an additional tablet device as the economy moves from Point A to Point B, and then also from Point C to Point D. Additionally, identify the opportunity cost of an additional smartphone as the economy moves from Point E to Point D, and then also from Point C to Point B. Explain whether your observations are consistent with the shape of this PPF.
Answer #1

Opportunity cost of additional tablets from A to B=

Decrease in smartphones / increase in tablets

= (55-50)/(25-15)= 0.5 units of smartphones.

Opportunity cost from point C to D= (40-25)/(45-35)

= 1.5 units of smartphones.

Opportunity cost of additional smartphones from E to D

= Decrease in tablets / increase in smartphones

= 10/25= 0.4 units of tablets.

Opportunity cost from C to B= 10/10= 1 unit of tablets

The observations are consistent with the shape of PPF.

The shape of PPF says that opportunity cost increases as we move to the right of the PPF. This makes it bowed outward.